Find the slope, x-intercept, and y-intercept of C=900+0.75y


1
Expert's answer
2021-11-07T19:46:32-0500

The slope is 0.75.

x-intercept is C = 900 + 0.75×0 = 900.

y-intercept is:

900 + 0.75y = 0,

y = -900/0.75 = -1200.